Compute the mass of the earth, assuming it to be a sphere of radius 6.37 x 10^6 m.


1
Expert's answer
2021-12-12T16:40:11-0500

Let's first find the volume of Earth:


"V=\dfrac{4}{3}\pi r^3."

Finally, we can find the mass of Earth from the definition of the density:


"\rho=\dfrac{m}{V},""m=\rho V=\dfrac{4}{3}\pi\rho r^3,""m=\dfrac{4}{3}\times\pi\times5510\ \dfrac{kg}{m^3}\times(6.37\times10^6\ m)^3=5.97\times10^{24}\ kg."